# Impact of Data Distribution on Fairness Guarantees in Equitable Deep Learning

### v1

@@ 3 Main Results
For a minimization problem with objective function $\Delta(f)$ and optimal solution $f^{*}=\arg\min_{f\in\mathcal{F}}\Delta(f)$, we say that $\hat{f}\in\mathcal{F}$ is an $\epsilon$-optimal solution if:
@@
where $\epsilon>0$ is the optimality gap. In the context of the fairness problem (Definition 3.1), this means:
$\max_{i,j}|E_{(x,y)\sim D_{a_{i}}}[\ell(\hat{f}(x),y)]-E_{(x,y)\sim D_{a_{j}}}[\ell(\hat{f}(x),y)]|\leq$
$\min_{f\in\mathcal{F}}\max_{i,j}|E_{(x,y)\sim D_{a_{i}}}[\ell(f(x),y)]-E_{(x,y)\sim D_{a_{j}}}[\ell(f(x),y)]|+\epsilon$
@@
The lemma follows from the standard uniform convergence bound for finite VC dimension function classes, applied to the pairwise differences of the loss functions for each pair of demographic groups. Now, consider the following algorithm:
  1. 1.
    Draw a sample of size $n\geq\frac{8M^{2}k^{2}}{\epsilon^{2}}(d\log(16Mk/\epsilon)+\log(4k^{2}/\delta))$.
  2. 2.
    For each $f\in\mathcal{F}$, calculate the empirical fairness error $\hat{\Delta}(f)=\max_{i,j}|\hat{L}_{i}(f)-\hat{L}_{j}(f)|$.
  3. 3.
    Return $\hat{f}=\arg\min_{f\in\mathcal{F}}\hat{\Delta}(f)$.
@@
$\Delta(\hat{f})$ $\leq\hat{\Delta}(\hat{f})+\epsilon/2$
$\leq\hat{\Delta}(f^{*})+\epsilon/2$
$\leq\Delta(f^{*})+\epsilon$
@@
where $D$ is the Lipschitz constant of functions in $\mathcal{F}$. Therefore, functions in $\mathcal{G}$ are bounded by $M$ and Lipschitz continuous with constant $LD$. By McDiarmid’s inequality, for any $f\in\mathcal{F}$, with probability at least $21-\delta/$ 2,
@@
By the union bound and the covering number bound, with probability at least $21-\delta/$ 2, for all $f\in\mathcal{F}$,
@@
$R(\hat{f}_{S})$ $\leq R_{\text{emp}}(\hat{f}_{S},S)+\epsilon$ $\leq R_{\text{emp}}(f^{,}S)+\epsilon$ $\leq R(f^{)}+2\epsilon$
@@
Lemma 3.16 (Risk Bound for Normal Distributions).
Let $p_{i}$ and $p$ be the density functions of two normal distributions with means $\mu_{i}$ and $\mu$, and covariance matrices $\Sigma_{i}$ and $\Sigma$, respectively. Then, for any function $f$:
$\left|\mathbb{E}_{(x,y)\sim p_{i}}[f(x,y)]-\mathbb{E}_{(x,y)\sim p}[f(x,y)]\right|\leq$ $L_{f}|\mu_{i}-\mu|_{2}+$
$L_{f}\sqrt{|\Sigma_{i}-\Sigma|_{F}}$
where $L_{f}$ is the Lipschitz constant of $f$ with respect to the Euclidean norm.
@@
Let $\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R})$ be the set of all probability measures on $\mathbb{R}^{d}\times\mathbb{R}$, and let $\Pi(p_{i},p)$ be the set of all joint probability measures on $(\mathbb{R}^{d}\times\mathbb{R})\times(\mathbb{R}^{d}\times\mathbb{R})$ with marginals $p_{i}$ and $p$. The Kantorovich-Rubinstein duality states that for any Lipschitz function $f$ with Lipschitz constant $L_{f}$, we have:
$\left|\mathbb{E}_{(x,y)\sim p_{i}}[f(x,y)]-\mathbb{E}_{(x,y)\sim p}[f(x,y)]\right|\leq L_{f}\cdot W_{1}(p_{i},p)$ (3)
where $W_{1}(p_{i},p)$ is the 1-Wasserstein distance between $p_{i}$ and $p$, defined as:
$W_{1}(p_{i},p)=\inf_{\gamma\in\Pi(p_{i},p)}\mathbb{E}_{((x,y),(x^{\prime},y^{\prime}))\sim\gamma}[|(x,y)-(x^{\prime},y^{\prime})|_{2}]$ (4)
Now, let’s focus on bounding the 1-Wasserstein distance between two normal distributions $p_{i}=\mathcal{N}(\mu_{i},\Sigma_{i})$ and $p=\mathcal{N}(\mu,\Sigma)$. By the triangular inequality, we have:
$W_{1}(p_{i},p)$ $\leq\inf_{\gamma\in\Pi(p_{i},p)}\mathbb{E}_{((x,y),(x^{\prime},y^{\prime}))\sim\gamma}[|x-x^{\prime}|2]+$ (5)
$\quad\inf_{\gamma\in\Pi(p_{i},p)}\mathbb{E}_{((x,y),(x^{\prime},y^{\prime}))\sim\gamma}[|y-y^{\prime}|_{2}]$ (6)
$\leq W_{1}(\mathcal{N}(\mu_{i},\Sigma_{i}),\mathcal{N}(\mu,\Sigma_{i}))+W_{1}(\mathcal{N}(\mu,\Sigma_{i}),\mathcal{N}(\mu,\Sigma))$ (7)
The first term in Eq. (7) can be bounded by the mean difference:
$W_{1}(\mathcal{N}(\mu_{i},\Sigma_{i}),\mathcal{N}(\mu,\Sigma_{i}))\leq|\mu_{i}-\mu|_{2}$ (8)
The second term in Eq. (7) can be bounded by the covariance difference [41]:
$W_{1}(\mathcal{N}(\mu,\Sigma_{i}),\mathcal{N}(\mu,\Sigma))\leq\sqrt{|\Sigma_{i}-\Sigma|_{F}}$ (9)
Combining Eq. (3), Eq. (7), Eq. (8), and Eq. (9), we obtain:
$\left|\mathbb{E}_{(x,y)\sim p_{i}}[f(x,y)]-\mathbb{E}_{(x,y)\sim p}[f(x,y)]\right|\leq$ $L_{f}|\mu_{i}-\mu|_{2}+$
$L_{f}\sqrt{|\Sigma_{i}-\Sigma|_{F}}$
which completes the proof. ∎
@@
Theorem 3.17 (Group-Specific Risk Bound Theorem for Normal Distributions).
@@
Let $\mathcal{F}$ be a function space with VC dimension $d_{\text{VC}}(\mathcal{F})$ and let $f^{*}_{i}=\arg\min_{f\in\mathcal{F}}R_{i}(f)$ be the risk minimizer for demographic group $a_{i}$. Let $\hat{f}_{S}=\arg\min_{f\in\mathcal{F}}R_{\text{emp}}(f,S)$ be the empirical risk minimizer based on a training set $S$ of size $m$ drawn from the overall data distribution. Then, under Assumptions 3.6 and 3.15, for any $\delta>0$, with probability at least $1-\delta$ over the random choice of $S$:
$R_{i}(\hat{f}_{S})-R_{i}(f^{*}_{i})\leq\frac{2LM}{\sqrt{m}}\left(\sqrt{2d{\text{VC}}(\mathcal{F})\ln\frac{em}{d{\text{VC}}(\mathcal{F})}}+\sqrt{2\ln\frac{4}{\delta}}\right)$
$\quad+L|\mu_{i}-\mu|_{2}+L\sqrt{|\Sigma_{i}-\Sigma|F}$
where $R_{i}(f)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[\ell(f(x),y)]$ is the expected risk for group $a_{i}$, $R_{\text{emp}}(f,S)=\frac{1}{m}\sum_{j=1}^{m}\ell(f(x_{j}),y_{j})$ is the empirical risk based on the training set $S$, $\mu$ and $\Sigma$ are the mean and covariance matrix of the overall data distribution, and $|\cdot|_{F}$ denotes the Frobenius norm.
Remark. Theorem 3.17 establishes that for a specific demographic group $a_{i}$ with a normal data distribution, the excess risk of the empirical risk minimizer trained on the overall data distribution, $R_{i}(\hat{f}_{S})-R_{i}(f^{*}_{i})$, can be bounded by three terms: The first term is the same as in Theorem 3.14 and depends on the VC dimension of the function space, the Lipschitz constant of the loss function, the boundedness of the loss function, and the size of the training set $S$. The second term is proportional to the Euclidean distance between the means of the group-specific distribution and the overall distribution, $|\mu_{i}-\mu|_{2}$. The third term is proportional to the square root of the Frobenius norm of the difference between the covariance matrices of the group-specific distribution and the overall distribution, $\sqrt{|\Sigma_{i}-\Sigma|_{F}}$. The result implies that the accuracy of the empirical risk minimizer on group $a_{i}$ depends not only on the size of the training set and the complexity of the function space but also on how close the group-specific distribution is to the overall distribution in terms of their means and covariance matrices. If the means and covariances are similar, the excess risk will be smaller, indicating better accuracy on group $a_{i}$. Conversely, if the means and covariances differ significantly, the excess risk will be larger, indicating potential accuracy disparities for group $a_{i}$.
The first term depends on the VC dimension of the hypothesis space, the Lipschitz constant of the loss function, the boundedness of the loss function, and the size of the training set $S$. The second term is proportional to the Euclidean distance between the means of the group-specific distribution and the overall distribution, $|\mu_{i}-\mu|_{2}$. The third term is proportional to the square root of the Frobenius norm of the difference between the covariance matrices of the group-specific distribution and the overall distribution, $\sqrt{|\Sigma_{i}-\Sigma|_{F}}$.
The result implies that the accuracy of the empirical risk minimizer on group $a_{i}$ depends not only on the size of the training set and the complexity of the hypothesis space but also on how close the group-specific distribution is to the overall distribution in terms of their means and covariance matrices. If the means and covariances are similar, the excess risk will be smaller, indicating better accuracy on group $a_{i}$. Conversely, if the means and covariances differ significantly, the excess risk will be larger, indicating potential accuracy disparities for group $a_{i}$.
@@
Let $\mathcal{F}$ be a function space with VC dimension $d_{VC}(\mathcal{F})$, and let $f^{*}=\arg\min_{f\in\mathcal{F}}R(f)$ and $f_{i}^{*}=\arg\min_{f\in\mathcal{F}}R_{i}(f)$ be the fairness and accuracy risk minimizers, respectively. Let $\hat{f}_{S}=\arg\min_{f\in\mathcal{F}}R_{emp}(f,S)$ be the empirical fairness risk minimizer based on a training set $S$ of size $m$ drawn from the overall data distribution. Then, under Assumptions 3.3, 3.6, and 3.15, for any $\delta>0$, with probability at least $1-\delta$ over the random choice of $S$:
$R_{i}(f^{*})-R_{i}(f_{i}^{*})\leq$ $\frac{4LM}{\sqrt{m}}\sqrt{2d_{VC}(\mathcal{F})\ln\frac{em}{d_{VC}(\mathcal{F})}+2\ln\frac{4}{\delta}}+$
$L|\mu_{i}-\mu|_{2}+L\sqrt{|\Sigma_{i}-\Sigma|_{F}}$
@@
By the triangle inequality and Theorems 3.14 and 3.17, we have:
$R_{i}(f^{*})-R_{i}(f_{i}^{*})$ $\leq R_{i}(f^{*})-R_{i}(\hat{f}_{S})+R_{i}(\hat{f}_{S})-R_{i}(f_{i}^{*})$
$\leq\frac{2LM}{\sqrt{m}}\sqrt{2d_{VC}(\mathcal{F})\ln\frac{em}{d_{VC}(\mathcal{F})}+2\ln\frac{4}{\delta}}$
$\quad+\frac{2LM}{\sqrt{m}}\sqrt{2d_{VC}(\mathcal{F})\ln\frac{em}{d_{VC}(\mathcal{F})}+2\ln\frac{4}{\delta}}+$
$\quad L|\mu_{i}-\mu|2+L\sqrt{|\Sigma_{i}-\Sigma|F}$
$=\frac{4LM}{\sqrt{m}}\sqrt{2d{VC}(\mathcal{F})\ln\frac{em}{d{VC}(\mathcal{F})}+2\ln\frac{4}{\delta}}+$
$\quad L|\mu_{i}-\mu|_{2}+L\sqrt{|\Sigma_{i}-\Sigma|_{F}}$
@@
Remark. Corollary 3.18 quantifies the trade-off between fairness and accuracy in equitable deep learning models. The left-hand side of the inequality, $R_{i}(f^{*})-R_{i}(f_{i}^{*})$, represents the difference in accuracy between the fairness risk minimizer $f^{*}$ and the accuracy risk minimizer $f_{i}^{*}$ for a specific demographic group $a_{i}$. The right-hand side provides an upper bound on this difference, which depends on the VC dimension of the function space, the Lipschitz constant of the loss function, the sample size, and the dissimilarity between the group-specific distribution and the overall distribution. This result has important implications for the design and evaluation of fair learning algorithms. It suggests that achieving perfect fairness (i.e., $R_{i}(f^{*})=R_{i}(f_{i}^{*})$ for all groups) may come at the cost of reduced accuracy, especially when the data distributions differ significantly across demographic groups. Practitioners should carefully consider this trade-off when developing equitable models and assess the impact of fairness constraints on the model’s performance for each group [19]. Also, the corollary highlights the importance of collecting representative data from each demographic group to mitigate the accuracy disparities induced by distributional differences. By reducing the dissimilarity between the group-specific distributions and the overall distribution, one can tighten the fairness-accuracy trade-off and achieve more equitable performance across all groups [2].
Theorem 3.19 (Expected Loss Bound for Demographic Group with Normal Distribution).
Let $\mathcal{D}_{a_{i}}\sim\mathcal{N}(\mu_{i},\Sigma_{i})$ be the data distribution for demographic group $a_{i}$, and let $\mathcal{D}\sim\mathcal{N}(\mu,\Sigma)$ be the overall data distribution. Let $f(\cdot)$ be a function that maps input $x$ to predicted labels $\hat{y}$, and let $\ell$ be a loss function bounded by $B$, i.e., $|\ell(f(x),y)|\leq B$ for all $x,y$. Suppose we have a training set ${(x_{j},y_{j})}_{j=1}^{n}$ of size $n$ drawn independently from $\mathcal{D}$. Then, with probability at least $1-\delta$, the expected loss of $f(\cdot)$ on $\mathcal{D}_{a_{i}}$ is bounded as follows:
$\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[\ell(f(x),y)]\leq$ $\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]+$
$B|\mu_{i}-\mu|_{2}+B\sqrt{|\Sigma_{i}-\Sigma|_{F}}$
@@
We start by applying the reference theorem to the overall data distribution $\mathcal{D}$:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]\leq\frac{1}{n}\sum_{j=1}^{n}\ell(f(x_{j}),y_{j})+B\sqrt{\frac{2|\Sigma|\log(2/\delta)}{n}}$
Now, we use Lemma 3.16 to relate the expected loss on $\mathcal{D}_{a_{i}}$ to the expected loss on $\mathcal{D}$: Applying the lemma with $f(x,y)=\ell(f(x),y)$, which has Lipschitz constant $B$ since $\ell$ is bounded by $B$, we have:
$\left|\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[\ell(f(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]\right|\leq$ $B|\mu_{i}-\mu|_{2}+$
$B\sqrt{|\Sigma_{i}-\Sigma|_{F}}$
Rearranging the terms, we get:
$\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[\ell(f(x),y)]\leq\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]+$ $B|\mu_{i}-\mu|_{2}+$
$B\sqrt{|\Sigma_{i}-\Sigma|_{F}}$
This completes the proof. ∎
Remark. Theorem 3.19 provides an insight into the performance of the empirical risk minimizer on a specific demographic group when the group’s data distribution differs from the overall training distribution. The theorem shows that the excess risk of the empirical risk minimizer on group $a_{i}$ can be bounded by three terms: (1) the excess risk of the empirical risk minimizer on the overall distribution, (2) the difference in means between the group-specific distribution and the overall distribution, and (3) the difference in covariance matrices between the group-specific distribution and the overall distribution. This result has important implications for understanding the sources of performance disparities across different demographic groups. It suggests that the accuracy of a model on a specific group depends not only on its performance on the overall population but also on how well the group’s data distribution matches the overall training distribution. If the means and covariances of the group-specific distribution differ significantly from those of the overall distribution, the model may suffer from poor accuracy on that group [2]. The theorem assumes that the group-specific distribution follows a normal distribution, which is a common assumption in statistical modeling and analysis [42]. The use of the Euclidean distance between means and the Frobenius norm of the covariance matrix difference allows for a concise and interpretable bound on the excess risk. The insights from Theorem 3.19 can inform the development of fair and accurate models. They highlight the importance of collecting representative data from each demographic group and considering the differences in data distributions when training and evaluating models [43]. Additionally, the theorem provides a way to quantify the impact of data distribution mismatch on the model’s performance, which can guide the development of targeted data collection and model improvement strategies [44].

### v2

@@ 3 Main Results
Let $\Delta(f)=\max_{i,j}\left|\mathbb{E}_{(x,y)\sim D_{a_{i}}}[\ell(f(x),y)]-\mathbb{E}_{(x,y)\sim D_{a_{j}}}[\ell(f(x),y)]\right|$ denote the fairness error of a function $f\in\mathcal{F}$. Let $f^{*}$ be the optimal fairness minimizer as defined in Theorem 3.4, i.e.,
$f^{*}=\arg\min_{f\in\mathcal{F}}\Delta(f).$
We say that $\hat{f}\in\mathcal{F}$ is an $\epsilon$-optimal solution if:
@@
where $\epsilon>0$ is the optimality gap. The expanded version is
$\max_{i,j}\left|\mathbb{E}_{(x,y)\sim D_{a_{i}}}[\ell(\hat{f}(x),y)]-\mathbb{E}_{(x,y)\sim D_{a_{j}}}[\ell(\hat{f}(x),y)]\right|\leq$
$\min_{f\in\mathcal{F}}\max_{i,j}\left|\mathbb{E}_{(x,y)\sim D_{a_{i}}}[\ell(f(x),y)]-\mathbb{E}_{(x,y)\sim D_{a_{j}}}[\ell(f(x),y)]\right|+\epsilon.$
@@
The lemma follows from the standard uniform convergence bound for finite VC dimension function classes, applied to the pairwise differences of the loss functions for each pair of demographic groups. Now, consider the following algorithm: draw a sample of size
$n\geq\frac{8M^{2}k^{2}}{\epsilon^{2}}\left(d\log\left(\frac{16Mk}{\epsilon}\right)+\log\left(\frac{4k^{2}}{\delta}\right)\right),$
then, for each $f\in\mathcal{F}$, compute the empirical fairness error $\hat{\Delta}(f)=\max_{i,j}|\hat{L}_{i}(f)-\hat{L}_{j}(f)|$, and finally return $\hat{f}=\arg\min_{f\in\mathcal{F}}\hat{\Delta}(f)$.
@@
[…]
@@
where $D$ is the Lipschitz constant of functions in $\mathcal{F}$. Therefore, functions in $\mathcal{G}$ are bounded by $M$ and Lipschitz continuous with constant $LD$. By McDiarmid’s inequality, for any $f\in\mathcal{F}$, with probability at least $1-\delta/2$,
@@
By the union bound and the covering number bound, with probability at least $1-\delta/2$, for all $f\in\mathcal{F}$,
@@
$R(\hat{f}_{S})$ $\leq R_{\text{emp}}(\hat{f}_{S},S)+\epsilon\$$\leq R_{\text{emp}}(f^{,}S)+\epsilon\$$\leq R(f^{)}+2\epsilon$
@@
Lemma 3.16 (Risk Bound for a Mixture of Normal Distributions).
Let the overall data distribution $\mathcal{D}$ be a mixture of $k$ demographic group distributions, $\mathcal{D}=\sum_{j=1}^{k}\pi_{j}\mathcal{D}_{a_{j}}$, where $\mathcal{D}_{a_{j}}\sim\mathcal{N}(\mu_{j},\Sigma_{j})$ is the data distribution for group $a_{j}$ and $\pi_{j}$ is its proportion in the overall population. Let $f(x,y)$ be a function that is $L_{f}$-Lipschitz with respect to its inputs $(x,y)$. The difference in expectation between any component distribution $\mathcal{D}_{a_{i}}$ and the overall mixture distribution $\mathcal{D}$ is bounded as follows:
[…]
@@
The expectation over the mixture distribution $\mathcal{D}$ can be written as a weighted sum of the expectations over its components:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}[f(x,y)]=\sum_{j=1}^{k}\pi_{j}\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{j}}}[f(x,y)]$
Since $\sum_{j=1}^{k}\pi_{j}=1$, we can express the expectation for group $a_{i}$ as $\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[f(x,y)]=\sum_{j=1}^{k}\pi_{j}\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[f(x,y)]$. The absolute difference in expectations is then:
$|\mathbb{E}_{\mathcal{D}_{a_{i}}}[f]-\mathbb{E}_{\mathcal{D}}[f]|$ $=\left|\sum_{j=1}^{k}\pi_{j}\mathbb{E}_{\mathcal{D}_{a_{i}}}[f]-\sum_{j=1}^{k}\pi_{j}\mathbb{E}_{\mathcal{D}_{a_{j}}}[f]\right|$
$=\left|\sum_{j=1}^{k}\pi_{j}\left(\mathbb{E}_{\mathcal{D}_{a_{i}}}[f]-\mathbb{E}_{\mathcal{D}_{a_{j}}}[f]\right)\right|$
By the triangle inequality, this is bounded by:
$\leq\sum_{j=1}^{k}\pi_{j}\left|\mathbb{E}_{\mathcal{D}_{a_{i}}}[f]-\mathbb{E}_{\mathcal{D}_{a_{j}}}[f]\right|$
For each pair of normal distributions $\mathcal{D}_{a_{i}}\sim\mathcal{N}(\mu_{i},\Sigma_{i})$ and $\mathcal{D}_{a_{j}}\sim\mathcal{N}(\mu_{j},\Sigma_{j})$, the Kantorovich-Rubinstein duality states that $|\mathbb{E}_{\mathcal{D}_{a_{i}}}[f]-\mathbb{E}_{\mathcal{D}_{a_{j}}}[f]|\leq L_{f}W_{1}(\mathcal{D}_{a_{i}},\mathcal{D}_{a_{j}})$, where $W_{1}$ is the 1-Wasserstein distance. The distance between two normal distributions is bounded by the differences in their means and covariance matrices:
[…]
Substituting this into the sum yields the final bound:
[…]
This completes the proof. ∎
@@
Theorem 3.17 (Group-Specific Risk Bound Theorem for Normal Mixture Distributions).
@@
Let $\mathcal{F}$ be a function space with VC dimension $d_{VC}(\mathcal{F})$ and let $f_{i}^{\dagger}=\arg\min_{f\in\mathcal{F}}R_{i}(f)$ be the risk minimizer for demographic group $a_{i}$. Let $\hat{f}_{S}=\arg\min_{f\in\mathcal{F}}R_{emp}(f,S)$ be the empirical risk minimizer based on a training set $S$ of size $m$ drawn from the overall mixture distribution $\mathcal{D}=\sum_{j=1}^{k}\pi_{j}\mathcal{D}_{a_{j}}$. Then, under Assumptions 3.6 and 3.15, for any $\delta>0$, with probability at least $1-\delta$ over the random choice of S:
[…]
Proof.
We decompose the excess risk as follows:
$\begin{split}R_{i}(\hat{f}_{S})-R_{i}(f_{i}^{\dagger})&=\left(R_{i}(\hat{f}_{S})-R(\hat{f}_{S})\right)+\left(R(\hat{f}_{S})-R(f_{i}^{\dagger})\right)\\ &+\left(R(f_{i}^{\dagger})-R_{i}(f_{i}^{\dagger})\right)\end{split}$
The first and third terms represent the shift between the group-specific risk $R_{i}$ and the overall risk $R$. Applying Lemma 3.16 to both terms, we can bound their absolute values:
[…]
The middle term, $R(\hat{f}_{S})-R(f_{i}^{\dagger})$, can be bounded using standard uniform convergence arguments. Let $\epsilon_{m}=\frac{2LM}{\sqrt{m}}\left(\sqrt{2d_{VC}(\mathcal{F})\ln\frac{em}{d_{VC}(\mathcal{F})}}+\sqrt{2\ln\frac{4}{\delta}}\right)$. By Lemma 3.13, with probability at least $1-\delta$, we have $\sup_{f\in\mathcal{F}}|R(f)-R_{emp}(f,S)|\leq\epsilon_{m}$. Following the standard argument for ERM consistency:
$R(\hat{f}_{S})$ $\leq R_{emp}(\hat{f}_{S},S)+\epsilon_{m}$
$\leq R_{emp}(f_{i}^{\dagger},S)+\epsilon_{m}\quad(\text{by definition of }\hat{f}_{S})$
$\leq R(f_{i}^{\dagger})+\epsilon_{m}+\epsilon_{m}=R(f_{i}^{\dagger})+2\epsilon_{m}$
So, $R(\hat{f}_{S})-R(f_{i}^{\dagger})\leq 2\epsilon_{m}$. Combining all parts yields:
$R_{i}(\hat{f}_{S})-R_{i}(f_{i}^{\dagger})$ $\leq|R_{i}(\hat{f}_{S})-R(\hat{f}_{S})|+(R(\hat{f}_{S})-R(f_{i}^{\dagger}))$
$\,+|R(f_{i}^{\dagger})-R_{i}(f_{i}^{\dagger})|$
$\leq 2\epsilon_{m}+2\times(\text{distribution shift bound})$
Substituting the expressions gives the final result. This completes the proof. ∎
Remark. Theorem 3.17 provides a crucial group-specific risk bound that explicitly quantifies how the statistical properties of a demographic group’s data distribution directly impact its model performance. The bound decomposes the excess risk for group $a_{i}$ into two interpretable components: a standard statistical learning term that converges at a rate of $O(1/\sqrt{m})$ with sample size, and a distributional shift term proportional to the sum of distances (in mean $\|\mu_{i}-\mu_{j}\|_{2}$ and covariance $\|\Sigma_{i}-\Sigma_{j}\|_{F}$) between group $a_{i}$ and all other groups $a_{j}$. This formalizes the intuitive notion that a group whose feature distribution deviates significantly from the overall training population will suffer higher prediction error, even if the model is optimally trained on the aggregate data. The theorem thus provides a theoretical foundation for observed empirical disparities, particularly in contexts like racial categories where distributional differences are often pronounced, and underscores that achieving true equity requires either explicitly accounting for these distributional shifts in the learning objective or ensuring the training data itself is representative and balanced.
@@
Let $\mathcal{F}$ be a function space. Let $f^{\dagger}=\arg\min_{f\in\mathcal{F}}R(f)$ be the accuracy risk minimizer over the overall mixture distribution $\mathcal{D}$, and let $f_{i}^{\dagger}=\arg\min_{f\in\mathcal{F}}R_{i}(f)$ be the accuracy risk minimizer for a specific demographic group $a_{i}$. Under the assumptions of a mixture of normal distributions (Assumption 3.15) and a Lipschitz-continuous loss function (Assumption 3.6), the trade-off between overall accuracy and group-specific accuracy is bounded by:
[…]
where $L$ is the Lipschitz constant of the loss function.
@@
By definition, $f_{i}^{\dagger}$ minimizes the group-specific risk $R_{i}(f)$, so the term we wish to bound, $R_{i}(f^{\dagger})-R_{i}(f_{i}^{\dagger})$, is non-negative. We can decompose this term by adding and subtracting the overall risk $R(f)$:
$\begin{split}R_{i}(f^{\dagger})-R_{i}(f_{i}^{\dagger})&=\left(R_{i}(f^{\dagger})-R(f^{\dagger})\right)+\left(R(f^{\dagger})-R(f_{i}^{\dagger})\right)\\ &+\left(R(f_{i}^{\dagger})-R_{i}(f_{i}^{\dagger})\right)\end{split}$
By the definition of $f^{\dagger}$ as the minimizer of the overall risk $R(f)$, we know that $R(f^{\dagger})\leq R(f_{i}^{\dagger})$. This implies that the middle term, $\left(R(f^{\dagger})-R(f_{i}^{\dagger})\right)$, is less than or equal to zero. We can thus bound the expression by dropping this non-positive term:
$R_{i}(f^{\dagger})-R_{i}(f_{i}^{\dagger})\leq\left(R_{i}(f^{\dagger})-R(f^{\dagger})\right)+\left(R(f_{i}^{\dagger})-R_{i}(f_{i}^{\dagger})\right)$
This is equivalent to:
$R_{i}(f^{\dagger})-R_{i}(f_{i}^{\dagger})\leq|R_{i}(f^{\dagger})-R(f^{\dagger})|+|R_{i}(f_{i}^{\dagger})-R(f_{i}^{\dagger})|$
Both terms on the right-hand side represent the difference between the expected loss over the group-specific distribution $\mathcal{D}_{a_{i}}$ and the overall mixture distribution $\mathcal{D}$ for a given function ($f^{\dagger}$ and $f_{i}^{\dagger}$, respectively). We can apply Lemma 3.16 to bound each of these terms:
[…]
Applying this bound for both $g=f^{\dagger}$ and $g=f_{i}^{\dagger}$ and summing the results yields the final inequality:
[…]
@@
Theorem 3.19 (Expected Loss Bound for a Demographic Group in a Normal Mixture Distribution).
Let $\mathcal{D}_{a_{i}}\sim\mathcal{N}(\mu_{i},\Sigma_{i})$ be the data distribution for group $a_{i}$, and let the overall data distribution be the mixture $\mathcal{D}=\sum_{j=1}^{k}\pi_{j}\mathcal{D}_{a_{j}}$. Let $l$ be a loss function that is $L_{l}$-Lipschitz and bounded by $B$. Then the expected loss of a function $f(\cdot)$ on group $a_{i}$ is bounded as follows:
[…]
@@
The proof follows directly from Lemma 3.16. We can write the expected loss on group $a_{i}$ as:
$\begin{split}\mathbb{E}_{\mathcal{D}_{a_{i}}}[l(f(x),y)]&=\mathbb{E}_{\mathcal{D}}[l(f(x),y)]\\ &+\left(\mathbb{E}_{\mathcal{D}_{a_{i}}}[l(f(x),y)]-\mathbb{E}_{\mathcal{D}}[l(f(x),y)]\right)\end{split}$
By taking the inequality, we get:
$\begin{split}\mathbb{E}_{\mathcal{D}_{a_{i}}}[l(f(x),y)]&\leq\mathbb{E}_{\mathcal{D}}[l(f(x),y)]\\ &+\left|\mathbb{E}_{\mathcal{D}_{a_{i}}}[l(f(x),y)]-\mathbb{E}_{\mathcal{D}}[l(f(x),y)]\right|\end{split}$
Applying the bound from Lemma 3.16 to the second term on the right-hand side, where the function is the loss $l(f(x),y)$ with Lipschitz constant $L_{l}$, immediately gives the desired result. This completes the proof. ∎
Remark. Theorem 3.19 supplies the first closed-form fairness certificate that translates readily measurable group statistics into an interpretable upper bound on the excess loss, thereby closing the gap between abstract OOD theory and practical fairness diagnostics. Specifically, classic domain-adaptation bounds [44, 45, 46] rely on the $\mathcal{H}\Delta\mathcal{H}$-divergence (where $\mathcal{H}$ is a hypothesis class) or generic Wasserstein distance and do not isolate the first- and second-order moment mismatches ($\|\mu_{i}-\mu\|_{2}$ and $\|\Sigma_{i}-\Sigma\|_{F}$) that fairness auditors can compute without labels. More recent distributionally-robust guarantees [47, 48] guard against worst-case perturbations inside a $\chi^{2}$ or Wasserstein ball rather than against a priori fixed demographic strata.
While Definition 3.1 relies on a continuous loss function to facilitate gradient-based optimization, our framework extends to deployment settings where only discrete decisions (e.g., 0-1 classification) are observable. To adopt the conclusions in such settings, the continuous loss $l$ (e.g., Cross-Entropy) serves as a differentiable, Lipschitz-continuous surrogate that upper-bounds the discrete 0-1 error. Under this interpretation, the bounds derived in Theorems 3.17 and 3.19 provide conservative guarantees: minimizing the distributional shift in feature space minimizes the upper bound on discrete decision disparities. Furthermore, as established in Eq. (2), the expected loss difference is a weighted aggregate of group-specific error rates ($L_{+,a_{i}},L_{-,a_{i}}$); therefore, theoretical constraints on the expected loss translate directly to constraints on observable fairness metrics, such as Equal Opportunity or Equalized Odds—in black-box deployments.